OpenStudy (anonymous):

find the least positive angle measurement that is coterminal with -40 degrees a. 320degrees b. 170degrees c. 420degrees d.288degrees

ganeshie8 (ganeshie8):

you may ADD/SUBTRACT 360, any number of times, to get coterminal angles

ganeshie8 (ganeshie8):

-40 = -40+360 = -40 + 360 + 360

OpenStudy (anonymous):

wait huh?!?! @ganeshie8

ganeshie8 (ganeshie8):

add 360 to the given angle, to get a positive coterminal angle

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):